# Predictions for Detecting a Turndown in the Baryonic Tully Fisher Relation

### v1

@@ 4.1 Non-circular motions and the bTFR
Although our spatially resolved bTFR slopes are generally consistent with observations (Table 1 and Section 3.3), we consider factors which could affect our rotation curve analysis.
@@
This uncertainty is why dwarf galaxies are often selected to be well-behaved before being included in velocity studies. Previous modelers have suggested that a galaxy’s disk may be out of equilibrium if the H i line-of-sight velocity dispersion is km s-1 (Verbeke et al., 2017), or not well modeled if the disk height-to-radius ratio is (Sands et al., 2024) or inclination is (Read et al., 2016). However, it has been found that discarding galaxies that are not well-behaved leads to a steeper bTFR slope (e.g., Stark et al., 2009; Downing & Oman, 2023). As previously mentioned, we have chosen to use all of our galaxies despite their morphology, which may explain discrepancies between our bTFR results and that of observations. In addition to this, it is likely that our bTFR slope would change if we were able to measure the rotation speeds with non-circular motions fully taken into account. We already see this in the V bTFR result, where measuring the rotation speed with the midplane potential gives a lower velocity than using the enclosed mass, presumably due to increasing importance of non-circular motions, and gives a slope of 3.36 instead of 3.69.
@@ 4.2 Can we measure V in dwarf galaxies with future observations?
Furthermore, to measure V in dwarf galaxies with spatially resolved observations, we need sufficiently low H i surface density limits. Our results suggest that solely observing galaxies at low rotation speeds is not enough to observe a turndown in the bTFR if we use the 1 M⊙ pc-2 surface density limit. Instead, measuring the rotation of diffuse H i gas at low densities is crucial. Figure 10 demonstrates that better sensitivity with a limit M⊙pc-2 is needed to possibly measure the bTFR turndown with H i . Surveys like MHONGOOSE (de Blok et al., 2024) and FEASTS (Wang et al., 2025) focus on measuring the diffuse H i gas in the outskirts of galaxies, down to surface densities M⊙ pc-2. MHONGOOSE has a sample of 30 galaxies, and FEASTS is designed to study 118 galaxies333https://github.com/FEASTS/LVgal/wiki, both with M M⊙. Although MHONGOOSE and FEASTS have a limited sample of dwarf galaxies, V could be successfully measured. Going deeper with MIGHTEE-H i or surveys with the Square Kilometer Array - Mid (Swart et al., 2022) could also increase the sample of galaxies with kinematic information beyond the 1 M⊙ pc-2 limit. If future H i surveys measure V for more dwarf galaxies at deeper sensitivities and find a steeper bTFR slope (), this may confirm a turndown. Observing a turndown in the bTFR would provide evidence that indeed, the inefficient galaxy formation of lower-mass galaxies alters the scaling between an H i disk and dark matter halo, and would suggest that it is not the same in all mass regimes.
Lastly, from Figure 6, it is not clear that it is possible to measure V using unresolved H i linewidths, as even W shows little evidence for a turndown. Certainly the bTFR slope steepens as we move from a measurement at W to W (Table 1). Yet even with decent spectral resolution (2.6 km s-1) and no observational noise, V is not within error of W for our simulated galaxies at MM⊙. This result is in striking contrast to the general wisdom that unresolved linewidths will be necessary to obtain statistics at low masses. Instead, these results suggest we might have to rely on a handful of well-measured resolved velocities at low galaxy masses in order to probe for evidence of a bTFR turndown.
@@ 5 Conclusions
We analyze the bTFR for a sample of 66 simulated dwarf galaxies from the Marvel-ous and Marvelous Massive Dwarfs simulations. These galaxies span a mass range of M M⊙. They are are primarily gas-dominated (Figure 1) and follow observed scaling relations for H i gas content and velocity widths (Figures 4 and 5).
@@
Our H i column density profiles predict that more gas can be observed past the fiducial limit of 1 M⊙ pc-2 (Figure 8), though more observations are needed to confirm the extended H i at radii beyond the 0.01 M⊙pc-2 limit (Figure 9). Better sensitivity is necessary to observe the more diffuse gas. We predict that a turndown in the bTFR (detectable as a slope ) could be measured using V if an H i sensitivity limit of 0.08 M⊙ pc-2 is achieved (Figure 10). This sensitivity is within reach of measurements in MIGHTEE-HI, MHONGOOSE, and FEASTS, though the number of observed dwarf galaxies that reach this depth may be small overall.
It is notable that our W unresolved H i linewidths are not sensitive to a bTFR turndown, and instead continue the power-law trend from higher mass galaxies. A steeper slope might be recoverable if W is measured, but this requires incredibly deep integrations with high signal-to-noise (Sardone et al., 2024), and will be difficult to achieve for more dwarf galaxies. It is generally assumed that unresolved H i linewidths must be used in order to obtain statistical samples of H i velocities at the dwarf scale. However, our results suggest that fewer dwarfs but with resolved rotation curves to lower H i surface densities might be the only way to observationally recover a turndown in the bTFR.

### v2

@@ 4 Discussion
4.1 Comparing bTFR Slopes
For the spatially resolved velocity methods, our bTFR slope using V () is slightly steeper than studies in a similar baryonic mass range, while the slope using V () is more consistent with observations. Lelli et al. (2019) found a bTFR slope of , and Šiljeg et al. (2024) found that classical dwarfs observed with Apertif followed a similar slope. Deg et al. (2024) used early WALLABY observations and found a slope of . Our slopes are less steep than the slope from Stark et al. (2009) ( (random) (systematic)), and this is likely due to their selection of gas-rich galaxies that reach a flat rotation curve, which has been shown to produce a slope closer to 4 (Bradford et al., 2016).
Dwarf galaxies exhibit more non-circular orbits due to baryonic feedback. With more dispersion-supported galaxies in the lower mass range, the discrepancy between assuming the mass-enclosed potential (V) versus the midplane gravitational potential (V) becomes larger. This may explain the steeper bTFR slope for V () versus V (). Non-circular motions are discussed further in Section 4.2, and will be explored in a future work.
For the spatially unresolved methods, our bTFR slope for W () is shallower, while the slopes for W () and W () are closer to observations. Using W, Bradford et al. (2016) measured a slope of , and Lelli et al. (2019) measured . We expect W and W to have steeper bTFR slopes than W, since W < W < W based on the shape of H i profiles, which become more Gaussian as mass decreases. This is consistent with previous studies (Bradford et al., 2016; El-Badry et al., 2017; Lelli et al., 2019; Sardone et al., 2024). Our bTFR slope using W is consistent with Bradford et al. (2016) and may differ from other studies due to how we treat inclination angle (as discussed in Section 2.3) and the fact that we include more irregular, lower-mass galaxies in our sample.
@@ 4.2 Non-circular motions and the bTFR
Although our spatially resolved bTFR slopes are generally consistent with observations (Section 4.1), we consider factors which could affect our rotation curve analysis.
@@
This uncertainty is why dwarf galaxies are often selected to be well-behaved before being included in velocity studies. Previous modelers have suggested that a galaxy’s disk may be out of equilibrium if the H i line-of-sight velocity dispersion is km s-1 (Verbeke et al., 2017), or not well modeled if the disk height-to-radius ratio is (Sands et al., 2024) or inclination is (Read et al., 2016). However, it has been found that discarding galaxies that are not well-behaved leads to a steeper bTFR slope (e.g., Stark et al., 2009; Downing & Oman, 2023). As previously mentioned, we have chosen to use all of our galaxies despite their morphology, which may explain discrepancies between our bTFR results and that of observations. In addition to this, it is likely that our bTFR slopes would change if we were able to measure the rotation speeds with non-circular motions fully taken into account. We already see this in the V bTFR result, where measuring the rotation speed with the midplane potential gives a lower velocity than using the enclosed mass, presumably due to increasing importance of non-circular motions, and gives a slope of 3.39 instead of 3.78.
@@ 4.3 Can we measure V in dwarf galaxies with future observations?
Furthermore, to measure V in dwarf galaxies with spatially resolved observations, we need sufficiently low H i surface density limits. Our results suggest that solely observing galaxies at low rotation speeds is not enough to observe a turndown in the bTFR if we use the 1 M⊙ pc-2 surface density limit. Instead, measuring the rotation of diffuse H i gas at low densities is crucial. Figure 10 demonstrates that better sensitivity with a limit M⊙pc-2 is needed to possibly measure the bTFR turndown with H i . Surveys like MHONGOOSE (de Blok et al., 2024) and FEASTS (Wang et al., 2025) focus on measuring the diffuse H i gas in the outskirts of galaxies, down to surface densities M⊙ pc-2. MHONGOOSE has a sample of 30 galaxies, and FEASTS is designed to study 118 galaxies333https://github.com/FEASTS/LVgal/wiki, both with M M⊙. Although MHONGOOSE and FEASTS have a limited sample of dwarf galaxies, V could be successfully measured. Going deeper with MIGHTEE-H i or surveys with the Square Kilometer Array - Mid (Swart et al., 2022) could also increase the sample of galaxies with kinematic information beyond the 1 M⊙ pc-2 limit. If future H i surveys measure V for more dwarf galaxies at deeper sensitivities and find a steeper bTFR slope (), this may confirm a turndown. Observing a turndown in the bTFR would provide evidence that indeed, the inefficient galaxy formation of lower-mass galaxies alters the scaling between an H i disk and dark matter halo, and would suggest that this scaling is not the same in all mass regimes.
Lastly, from Figure 6, it is not clear that it is possible to measure V using unresolved H i linewidths, as even W shows little evidence for a turndown. Certainly the bTFR slope steepens as we move from a measurement at W to W ( to ). This work focuses on how different velocity definitions impact the bTFR, and therefore we have not performed mock observations or beam smearing effects. Yet even with these ‘ideal’ H i profiles, with decent spectral resolution (2.6 km s-1) and no observational noise, V is not within error of W for our simulated galaxies at MM⊙. This result is in striking contrast to the general wisdom that unresolved linewidths will be necessary to obtain statistics at low masses. Instead, these results suggest we might have to rely on a handful of well-measured resolved velocities at low galaxy masses in order to probe for evidence of a bTFR turndown.
@@ 5 Conclusions
We analyze the bTFR for a sample of 75 simulated dwarf galaxies from the Marvel-ous and Marvelous Massive Dwarfs simulations. These galaxies span a mass range of M M⊙. They are are primarily gas-dominated (Figure 1) and follow observed scaling relations for H i gas content and velocity widths (Figures 4 and 5).
@@
Our H i column density profiles predict that more gas can be observed past the fiducial limit of 1 M⊙ pc-2 (Figure 8), though more observations are needed to confirm the extended H i at radii beyond the 0.01 M⊙pc-2 limit (Figure 9). Better sensitivity is necessary to observe the more diffuse gas. We predict that a turndown in the bTFR within CDM (detectable as a slope ) could be measured using V if an H i sensitivity limit of 0.08 M⊙ pc-2 is achieved (Figure 10). This sensitivity is within reach of measurements in MIGHTEE-HI, MHONGOOSE, and FEASTS, though the number of observed dwarf galaxies that reach this depth may be small overall.
It is notable that our W unresolved H i linewidths are not sensitive to a bTFR turndown, and instead continue the power-law trend from higher mass galaxies. A steeper slope might be recoverable if W is measured, but this requires incredibly deep integrations with high signal-to-noise (Sardone et al., 2024), and will be difficult to achieve for more dwarf galaxies. It is generally assumed that unresolved H i linewidths must be used in order to obtain statistical samples of H i velocities at the dwarf scale. However, our results suggest that fewer dwarfs with resolved rotation curves to lower H i surface densities might be the only way to observationally recover a turndown in the bTFR.
@@ Acknowledgements
Data Availability Statement
Data from the Marvel-ous and Massive Dwarfs simulations have not been publicly released, but can be provided upon request to the author(s).
Appendix A H i Profiles for the Full Sample
The full sample of 75 dwarf galaxies and their H i emission profiles are shown in Figure 11. The galaxies are ordered by their total H i mass and span a range of MM. Galaxies with lower masses tend to have more Gaussian H i profiles, while higher mass galaxies exhibit the double-horn feature with high-velocity gas rotating in the outermost regions of the disk. Despite the criteria for these galaxies to be isolated (no galaxy of equal or greater mass within the virial radius) and in the field, some of the H i profiles are asymmetric, which will be explored in future work.
Appendix B Estimate of Observational Errors
The bTFR compares a galaxy’s baryonic mass (MM⋆+1.4M) and rotation velocity. Here, we elaborate on our estimate of observational errors for each axis. We do not perform mock observations for our simulated dwarf galaxies. However, we attempt to plot our bTFR trends with reasonable errors to determine the significance of differences between H i and halo velocities.
Our baryonic mass error must account for errors in stellar mass and H i mass. For stellar mass, we use the mean uncertainty from Bradford et al. (2015) of 0.164 dex. This value is chosen since their sample covers a similar mass range with M M⋆. For H i mass errors, we adopt the error function from Haynes et al. (2018), which propagates distance () and flux () percent uncertainties:
For this error estimate, we use a signal-to-noise ratio of 10 since this is the threshold used by Sardone et al. (2024) to ensure robust measurements of W and W. We adopt a distance uncertainty of , which is the average for the SPARC (Spitzer Photometry and Accurate Rotation Curves) galaxies (Lelli et al., 2016a). Based on this distance uncertainty, we find an H i mass uncertainty of 0.161 dex. Since stellar masses and H i masses are typically measured independently through different methods, we assume Gaussian statistics and take the log-space stellar and H i mass errors in quadrature to get 0.23 dex as the baryonic mass error. Gaussian statistics are commonly adopted for error propagation in bTFR studies (e.g., Stark et al., 2009; Lelli et al., 2016b; Papastergis et al., 2016; McQuinn et al., 2022).
For the total error in velocity, we consider the uncertainty from turbulent motion (dispersion) and the uncertainty in measuring the rotation speed. The asymmetric drift correction from gas turbulence is usually approximated with a value of km s-1 (Stilp et al., 2013b). By taking the logarithm of the upper limit as a fraction of the fiducial value, we get dex (25% error) as the dispersion uncertainty. For the uncertainty of spatially resolved velocities, we use the average error of from McQuinn et al. (2022), in which their sample had lower-mass galaxies with M M⊙. McQuinn et al. (2022) similarly assumed Gaussian statistics and added the errors for dispersion and rotation speed in quadrature (Equation 2 of their work), and we find 0.11 dex (29%) for the velocity error (Figure 6, top panels). For the uncertainty of H i linewidths, we use 9%, which is the average from Bradford et al. (2016). Adding the errors in quadrature gives a total velocity uncertainty of 0.10 dex (27% error), as shown in the bottom row of Figure 6.
Appendix C Relations between Velocity Methods
We determine best-fit relationships between the H i velocities and the maximum rotation speeds, V and V. Table 1 summarizes the best-fit linear relations between the H i velocity methods to the halos speeds V (which uses the mass-enclosed gravitational potential) and V (which uses the midplane gravitational potential). In this table, we also include the best-fit linear relation between maximum halo speeds V and V.
| y | m | x | b |
| V | V | ||
| V | V | ||
| W | V | ||
| W | V | ||
| W | V | ||
| V | V | ||
| V | V | ||
| W | V | ||
| W | V | ||
| W | V | ||
| Vmax,mid | Vmax,circ |